OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is 6, the last term is 24,576, and the common ratio is -4? -32,766 -19,662 19,662 32,766

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

i WILL MEDAL/FAN PLEASE HELP!

OpenStudy (alekos):

\[\sum_{i=1}^{n}a _{i} = a _{1}\frac{ 1-r ^{n} }{1-r }\]

OpenStudy (anonymous):

I dont understand the formula

OpenStudy (alekos):

just plug the values in the RHS and you'll get the sum

OpenStudy (alekos):

a1=6, r=-4 and n=7

OpenStudy (anonymous):

What values?

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

I got 19,664.4

OpenStudy (anonymous):

ans is -19662

OpenStudy (anonymous):

How did u get that?

OpenStudy (anonymous):

@AHSHAJHASJDHS ur ans is wrong,,,,according to me...

OpenStudy (alekos):

no your both wrong

OpenStudy (anonymous):

ok wait i am expplaining

OpenStudy (anonymous):

Can you help me get to the right answer? Mine didnt even match

OpenStudy (alekos):

\[6 \frac{ 1-(-4)^{7} }{1-(-4)}\]

OpenStudy (alekos):

this is what you need to calculate

OpenStudy (anonymous):

I got 19,662

OpenStudy (anonymous):

(6) 16385 ------- = 3277 (6) = 19662 5

OpenStudy (anonymous):

Is that right?

OpenStudy (alekos):

Yep, thats it